### v2

# SpiralMLP: A Lightweight Vision MLP Architecture

Abstract
We present SpiralMLP, a novel architecture introduces a Spiral FC layer as a replacement for the conventional Token Mixing approach. Differing from several existing MLP-based models that primarily emphasize axes, our Spiral FC layer is designed as a deformable convolution layer with spiral-like offsets. We further adapt Spiral FC into two variants: Self-Spiral FC and Cross-Spiral FC, enabling both local and global feature integration seamlessly, eliminating the need for additional processing steps. To thoroughly investigate the effectiveness of the spiral-like offsets and validate our design, we conduct ablation studies and explore optimal configurations. In empirical tests, SpiralMLP reaches state-of-the-art performance, similar to Transformers, CNNs, and other MLPs, benchmarking on ImageNet-1k, COCO and ADE20K. SpiralMLP still maintains linear computational complexity $O(HW)$ and is compatible with varying input image resolutions. Our study reveals that targeting the full receptive field is not essential for achieving high performance, instead, adopting a refined approach offers better results.
1 Introduction
1.1 Background
Earlier image classification systems mainly relied on CNN-based architectures [20, 84, 52, 56], which excel with controlled datasets but struggle with biased or uncontrolled conditions. Subsequently, Transformer-based architectures [66, 14, 27, 2] have emerged as alternatives, benefiting from self-attention mechanism that excel with large datasets [53] and are adaptable for various tasks [42]. However, they are often more expensive in pretraining and need specific datasets for better performance on downstream tasks.
MLP-based architectures [59, 39] have also shown promise in computer vision tasks, matching Transformer performance with a more data-efficient and lighter design. These systems use two main components: Channel Mixing, which projects features along the channel dimension, and Token Mixing, which captures spatial information by projecting feature along the spatial dimension. These mixing layers collectively enhance context aggregation, improving robustness and reducing training resource needs.
Figure 1: (a) While the Channel FC concentrates solely at the target point, marked with a $\bigstar$, the Spiral FC captures richer spatial information. Spiral FC is in accordance with Eqs. 4 and 5, the input channel dimension $C_{\text{in}}=14$, the maximum amplitude $A_{\text{max}}=6$ and $T=8$. The coordinate numbers are arranged as $(H,W,C)$. This illustrative example only contains half of the $C_{\text{in}}$. (b) provides a complete visualization when the parameters are: $C_{\text{in}}=20$, $A_{\text{max}}=3$ and $T=8$.
Figure 2: (a) displays the comprehensive architecture of SpiralMLP in PVT-style, featuring four distinct stages. Each stage is composed of multiple Patch Embedding layers and identically-configured Spiral Blocks. (b) explores the internal layout of a Spiral Block, where the proposed Spiral Mixing replaces the traditional Token Mixing. (c) outlines the components of Spiral Mixing, which incorporates the meticulously designed Spiral FC to effectively capture spatial information.
1.2 MLP-Based Architectures.
The pioneering MLP-Mixer [59] proposes a simple yet powerful architecture with both Token Mixing and Channel Mixing. Given a feature map $X\in\mathbb{R}^{H\times W\times C_{\text{in}}}$, where $H$, $W$ are the height and weight, $C_{\text{in}}$ is the input channel dimension, let $W^{\text{Tmix}}\in\mathbb{R}^{H\cdot W\times H\cdot W}$ denote the token mixing weight matrix, the operation applied to the reshaped input $X^{T}\in\mathbb{R}^{C_{\text{in}}\times H\cdot W}$ is described as follows:
$\text{Tmix}(X)=(X^{T}W^{\text{Tmix}})^{T}$ (1)
where, $\mathbb{R}^{H\cdot W}$ indicates the dimensions are flattened while $\mathbb{R}^{H\times W}$ denotes the dimensions are separated, and $\text{Tmix}(\cdot)\in\mathbb{R}^{H\cdot W\times C_{\text{in}}}$ is the output of token mixing. Eq. 1 is to simulate the attention operation to integrate spatial information, it is followed by the channel mixing that operates along the channel dimension. We define the channel mixing weight matrix as $W^{\text{Cmix}}\in\mathbb{R}^{C_{\text{in}}\times C_{\text{out}}}$, the channel mixing output $\text{Cmix}(\cdot)\in\mathbb{R}^{H\times W\times C_{\text{out}}}$ is expressed as follows:
$\text{Cmix}(\text{Tmix}(X))=(X^{T}W^{\text{Tmix}})^{T}W^{\text{Cmix}}$ (2)
While MLP-Mixer shows strong performance, it is limited by its quadratic computational complexity $O(H^{2}W^{2})$ (Eq. 1) and requires fixed image sizes due to its fully-connected token mixing layer. Alternatives like gMLP [39] introduces a spatial gating unit for better integration, FNet [32] uses Fourier transforms for token mixing, and HireMLP [18] mimics self-attention by swapping elements across regions. Other developments include ResMLP [60], which replaces LayerNorm with trainable matrices, s2MLP [76] utilizes shifting operations, WaveMLP [58] treats pixels as complex numbers, ViP [23] employs a permutator for spatial data, and MorphMLP [79] gradually expands its receptive field.
Despite advancements, the MLPs mentioned earlier haven’t significantly cut computational complexity, leaving an opening for SparseMLP [57], ASMLP [36], CycleMLP [6], and ATM [69]. SparseMLP and ASMLP use dense token mixing along the channel dimension, while CycleMLP introduces Cycle FC for sparser, channel-wise mixing with fixed offsets $S_{H}$ and $S_{W}$. ATM [69], on the other hand, uses trainable offsets for dynamic token mixing. However, these models restrict token mixing to horizontal and vertical axes, limiting their ability to integrate feature information across different spatial dimensions.
To address these challenges, we present SpiralMLP with its core component, Spiral FC, based on Channel FC as shown in Fig. 1(a). Spiral FC offers sufficient receptive field coverage while maintaining linear computational complexity. The paper is organized as follows:
  • We introduce the SpiralMLP architecture and its foundational Spiral FC layer.
  • We conduct experiments to demonstrate SpiralMLP’s superiority over other state-of-the-art models.
  • We perform ablation studies to explore optimal configurations, followed by conclusions and discussions on potential future improvements.
2 Methodology
2.1 Spiral FC
We aim to design a compact token mixing layer that captures spatial information efficiently. Our review indicates that traditional designs with criss-cross fully-connected layers fail to optimize the offset function, resulting in inadequate spatial coverage. To address these challenges, we draw inspiration from natural spiral patterns and AttentionViz [75], noted for its spiral patterns in transformer attention visualizations.
As a result, we introduce the Spiral Fully-Connected Layer (Spiral FC), intended to replace standard Token Mixing (Eq. 1) in the MLP-Mixer architecture. Described in Fig. 1(b), Spiral FC leverages a spiral trajectory across the feature map $X\in\mathbb{R}^{H\times W\times C_{\text{in}}}$:
$\text{Spiral FC}_{i,j,:}(X)=\sum^{C_{\text{in}}}_{c=0}X_{i+\phi_{i}(c),j+\phi_{j}(c),c}W^{\text{spiral}}_{c,:}+b^{\text{spiral}}$ (3)
where, $W^{\text{spiral}}\in\mathbb{R}^{C_{\text{in}}\times C_{\text{out}}}$, $b^{\text{spiral}}\in\mathbb{R}^{C_{\text{out}}}$ are the trainable matrix and bias, $\text{Spiral FC}_{i,j,:}(\cdot)$ is the output at position $(i,j,:)$. Both $\phi_{i}(c)$ and $\phi_{j}(c)$ serve as the offset functions along $H,W$ axes respectively within $X$. Furthermore, with the central axis of the spiral trajectory aligns along the channel dimension, the offset functions $\phi_{i}(c)$ and $\phi_{j}(c)$ are defined in a spiral manner:
$\phi_{i}(c)=A(c)\cos{(\frac{c\times 2\pi}{T})}$ (4)
$\phi_{j}(c)=A(c)\sin{(\frac{c\times 2\pi}{T})}$ (5)
where, $T$ is the constant period, $A(\cdot)$ is the amplitude that controls the width of the spiral trajectory, for conciseness, we formulate the amplitude function $A(\cdot)$ with the basic pattern 111 Sec. 4.1 provides additional cases with universal offset functions.:
$A(c)=\begin{cases}\lfloor\frac{2A_{\text{max}}}{C_{\text{in}}}c\rfloor,&0\leq c<\frac{C_{\text{in}}}{2}\\ \lfloor 2A_{\text{max}}-\frac{2A_{\text{max}}}{C_{\text{in}}}c\rfloor,&\frac{C_{\text{in}}}{2}\leq c\leq C_{\text{in}}\end{cases}$ (6)
where, $A_{\text{max}}$ is the maximum amplitude. When $A_{\text{max}}=0$, the Spiral FC is identical to Channel FC, denoted as Self-Spiral FC. Conversely, when $A_{\text{max}}\neq 0$, it is termed as Cross-Spiral FC. Additionally, we employ a sliding window with a stepsize of $1$. It not only makes the Spiral FC agnostic to the input size, but also enables the flexible feature extraction through meticulously modifying the offset functions (Eqs. 4 and 5), thereby ensuring the Spiral FC operate with linear computational complexity.
2.2 Spiral Mixing
At a specific position $(i,j,:)$, Self-Spiral FC captures the local information from itself, yielding an output denoted as $X^{\text{self}}_{i,j,:}$. Conversely, Cross-Spiral FC selectively incorporates spatial information from within the receptive field which is determined by $A_{\text{max}}$, and the output is represented as $X^{\text{cross}}_{i,j,:}$. Across the whole feature map, both the Self-Spiral FC and Cross-Spiral FC operate in parallel, and their outputs, $X^{\text{self}}\in\mathbb{R}^{H\times W\times C_{\text{out}}}$ and $X^{\text{cross}}\in\mathbb{R}^{H\times W\times C_{\text{out}}}$, merge together in the subsequent Merge Head [^2] Detailed explanation is provided in Appendix.:
$a=\sigma(W^{\text{merge}}\times[\frac{1}{HW}\sum^{HW}_{i=0}\mathcal{F}(X^{\text{self}}+X^{\text{cross}})_{i,:}])$ (7)
where, the reshaping function $\mathcal{F}:\mathbb{R}^{H\times W\times C_{\text{out}}}\rightarrow\mathbb{R}^{H\cdot W\times C_{\text{out}}}$ flattens the first two dimensions of the input, creating a new projection along the $HW$ dimension. Then, the newly generated projection is averaged into $\mathbb{R}^{1\times C_{\text{out}}}$. Subsequently, $W^{\text{merge}}\in\mathbb{R}^{2,1}$ maps this average from $\mathbb{R}^{1\times C_{\text{out}}}$ to $\mathbb{R}^{2\times C_{\text{out}}}$. Finally, the SoftMax function $\sigma(\cdot)$ determines the weights $a\in\mathbb{R}^{2\times C_{\text{out}}}$. Then at position $(i,j,:)$, the Merge Head generates the output:
$X^{\text{spiral}}_{i,j,:}=a_{1,:}\odot X^{\text{self}}_{i,j,:}+a_{2,:}\odot X^{\text{cross}}_{i,j,:}$ (8)
where, $\odot$ represents the element-wise multiplication. The weights $a$ is to modulate the contribution of the inputs. Furthermore, across the entire $X^{\text{spiral}}$, the weights $a$ is broadcast to influence all elements in both $X^{\text{self}}$ and $X^{\text{cross}}$.
Collectively, Self-Spiral FC, Cross-Spiral FC and Merge Head together constitute the Spiral Mixing, as depicted in Fig. 2 (c). Spiral Mixing transforms the input feature map $X\in\mathbb{R}^{H\times W\times C_{\text{in}}}$ to $X^{\text{spiral}}\in\mathbb{R}^{H\times W\times C_{\text{out}}}$, functioning similarly to vanilla Token Mixing.
2.3 Spiral Block
The output $X^{\text{spiral}}$ of Spiral Mixing subsequently proceeds to the Channel Mixing structured as a MLP with a GeLU [21] activation function $\zeta(\cdot)$:
$X^{\text{chn}}=\zeta(X^{\text{spiral}}\times W^{\text{mlp1}})\times W^{\text{mlp2}}$ (9)
where, $W^{\text{mlp1}}\in\mathbb{R}^{C_{\text{out}}\times C_{\text{mlp}}}$ and $W^{\text{mlp2}}\in\mathbb{R}^{C_{\text{mlp}}\times C_{\text{out}}}$ are the linear layer weight matrices. $X^{\text{chn}}$ is the output of Channel Mixing.
Spiral Mixing and Channel Mixing collectively compose the Spiral Block, as depicted in Fig. 2 (b). To summarize, Spiral Block accepts the feature map $X\in\mathbb{R}^{H\times W\times C_{\text{in}}}$ as the input, and initially processes it through a LayerNorm [1] before the Spiral Mixing. Then it produces $X^{\prime}$ integrated with a residual connection. Following this, $X^{\prime}$ is processed through another LayerNorm and then Channel Mixing, coupled with another residual connection, resulting in the output $Y$:
$X^{\prime}$ $=\text{Spiral Mixing}(\text{LN}(X))+X$ (10)
$Y$ $=\text{Channel Mixing}(\text{LN}(X^{\prime}))+X^{\prime}$ (11)
2.4 Overall Architecture and Model Zoo
We firstly construct our SpiralMLP based on the PVT [68] framework, the models are scaled from SpiralMLP-B1 to SpiralMLP-B5 by adjusting the hyperparameters. In each model, 4 stages are integrated, and the spatial resolution is reduced while the channel dimension is increased along with the process. Thereby it facilitates effective down-sampling of spatial resolution and optimizes computational efficiency. A depiction of the PVT-style SpiralMLP architecture can be found in Fig. 2 (a).
Furthermore, we have also developed variants modeled after the Swin architecture. The models are categorized into three types: SpiralMLP-T (Tiny), SpiralMLP-S (Small), and SpiralMLP-B (Base). The structural details of both PVT-style and Swin-style will be further provided in the appendix.
3 Experiments
Model CIFAR-10(%) CIFAR-100(%) Params(M)
Spiral-B1 (ours) 95.6 78.6 14
CaiT [62] 94.9 76.9 9
MONet-T [7] 94.8 77.2 10.3
Cycle-B1 [6] 94.5 77.3 15
PiT [22] 94.2 75.0 7
Swin [42] 94.0 77.3 7
VGG19-bn [52] 94.0 72.2 39
ResNet50 [20] 93.7 77.4 24
ViT [14] 93.6 73.8 3
Swin-v2-T [41] 89.7 70.2 28
Table 1: Top-1 accuracy achieved through training from scratch on both CIFAR-10 and CIFAR-100.
Model Top-1 Params FLOPs Model Top-1 Params FLOPs
Acc (%) (M) (G) Acc (%) (M) (G)
SpiralMLP-B5 (ours) 84.0 68 11.0 Swin-B [42] 83.5 88 15.4
SpiralMLP-B4 (ours) 83.8 46 8.2 gSwin-S [17] 83.0 19 4.2
SpiralMLP-B (ours) 83.6 67 11.0 SimA-XCiT-S12/16 [28] 82.1 26 4.8
SpiraMLP-S (ours) 83.3 56 9.1 SimA-CvT-13 [28] 81.4 20 4.5
ATMNet-L [69] 83.8 76 12.3 SimA-DeiT-S [28] 79.8 22 4.6
HireMLP-Large [18] 83.8 96 13.4 NOAH [33] 77.3 26 -
WaveMLP-B [58] 83.6 63 10.2 CRATE-L [77] 71.3 78 -
MorphMLP-L [79] 83.4 76 12.5 CRATE-B [77] 70.8 23 -
MorphMLP-B [79] 83.2 58 10.2 DeepMAD-89M [51] 84.0 89 15.4
CycleMLP-B [6] 83.4 88 15.2 DeepMAD-50M [51] 83.9 50 8.7
CycleMLP-B5 [6] 83.1 76 12.3 EfficientNet-B4 [56] 82.6 19 4.2
sMLP-B [57] 83.4 66 14.0 VanillaNet-13-1.5 [4] 82.5 128 26.5
ASMLP-B [36] 83.3 88 15.2 VanillaNet-13 [4] 82.1 59 11.9
gMLP [39] 81.6 45 31.6 HGRN-DeiT-Small [47] 80.1 24 -
ConvMixer-1536/20 [65] 81.4 52 - HGRN-DeiT-Tiny [47] 74.4 6 -
ConvMixer-1536/20 [65] 80.4 49 - ResNet-50 [20] 75.3 25 3.8
MONet-S [7] 81.3 33 6.8 Vim-S [83] 80.5 26 -
MONet-T [7] 77.0 10 2.8 Vim-Ti [83] 78.3 7 -
ResMLP-B24 [60] 81.0 116 23.0 M2-ViT-b [15] 79.5 45 -
S2MLP-deep [76] 80.7 51 10.5 ViT-b-Monarch [15] 78.9 33 -
S2MLP-wide [76] 80.0 71 14.0 HyenaViT-b [46] 78.5 88 -
ConvMLP-L [34] 80.2 43 9.9 RepMLP-Res50-g8/8 [13] 76.4 59 12.7
ConvMLP-M [34] 79.0 17 4.0 MLPMixer-B/16 [59] 76.4 59 12.7
RepMLP-Res50-g4/8 [13] 80.1 87 8.2 AFFNet [25] 79.8 6 1.5
Table 2: Top-1 accuracy on ImageNet-1k, with $224\times 224$ as the input resolution. In terms of background colors, , , , denote MLPs, Transformers, CNNs and State-Space Models, respectively.
Figure 3: In contrast to ASMLP [36] and Swin [42], our SpiralMLP demonstrates superior object-focused attention. SpiralMLP exhibits enhanced sensitivity, especially for elongated or curved objects. The backbones employed for heatmaps generation are SpiralMLP-B5, ASMLP-B and Swin-B. The images are sourced from the ImageNet-1k [50] validation dataset, with corresponding labels.
We initially perform experiments with SpiralMLP-B1 on CIFAR-10 [29] and CIFAR-100 [29], comparing it against architectures of similar scale, including MLPs, CNNs, and Transformers. The outcomes are presented in Tab. 1, all of the models are trained from scratch.
We extend our experimentation to include image classification on ImageNet-1k [50], as well as object detection and instance segmentation on the COCO [38]. Furthermore, we assess its semantic segmentation capabilities on ADE20K [82].
3.1 Image Classification on ImageNet-1k
3.1.1 Settings
Our implementation primarily draws from DeiT [61]. The training is 4 NVIDIA A100 GPUs for a total of 300 epochs. The overall batch size is 512 and we employ the Top-1 accuracy for image classification.
3.1.2 Comparison with MLPs
As shown in Tab. 2 , SpiralMLP-B achieves a Top-1 accuracy of $84.0\%$ on the ImageNet-1k, with the input resolution of $224\times 224$. This performance notably exceeds that of the best-performing models of ATMNet-L [69], HireMLP-Large [18], WaveMLP-B [58], MorphMLP-L [79] and CycleMLP-B [6], by +$0.2\%$, +$0.2\%$, +$0.4\%$, +$0.6\%$ and +$0.6\%$, respectively. Furthermore, compared to the S2MLP-wide [76], which has a similar model size with $71$M parameters, SpiralMLP surpasses it by +$4.0\%$ with only $68$M parameters. In addition to the advantage on the model size, SpiralMLP also demonstrates potential balance between computational efficiency and accuracy. It is evident that among a cohort of models with accuracy exceeding $83\%$ (including ATMNet-L [69], HireMLP-Large [18], WaveMLP-B [58], MorphMLP-B [79], MorphMLP-L [79], CycleMLP-B [6], CycleMLP-B5 [6], sMLP-B [57] and ASMLP-B [36], SpiralMLP-B5 stands out due to a lower FLOPs of $11.0$G and the highest accuracy.
BackBone RetinaNet 1×1$\times$1 ×
Params (M) FLOPs (G) AP AP50 AP75 APS APM APL
ResNet101 [20] 56.7 492.2 38.5 57.8 41.2 21.4 42.6 51.1
ConvMLP-L [34] 52.9 - 40.2 59.3 43.3 23.5 43.8 53.3
ResNeXt101-64x4d [71] 95.5 - 41.0 60.9 44.0 23.9 45.2 54.0
CycleMLP-B5 [6] 85.9 360.3 42.7 63.3 45.3 24.1 46.3 57.4
ATMNet-L [69] 86.0 405.0 46.1 67.4 49.4 29.9 50.1 61.0
PVTv2-B5 [68] 91.7 - 46.2 67.1 49.5 28.5 50.0 62.5
SpiralMLP-B5 (ours) 79.8 325.0 46.5 67.7 49.8 30.3 50.8 62.8
BackBone Mask R-CNN 1×1$\times$1 ×
Params (M) FLOPs (G) AP AP50 AP75 APS APM APL
ResNet101 [20] 63.2 - 40.4 61.1 44.2 36.4 57.7 38.8
ConvMLP-L [34] 62.2 - 41.7 62.8 45.5 38.2 59.9 41.1
Swin-T [42] 47.8 267.0 42.7 65.2 46.8 39.3 62.2 42.2
ResNeXt101-64x4d [71] 101.9 - 42.8 63.8 47.3 38.4 60.6 41.3
VanillaNet-13 [4] 76.3 421.0 42.9 65.5 46.9 39.6 62.5 42.2
CycleMLP-B5 [6] 95.3 - 44.1 65.5 48.4 40.1 62.8 43.0
HireMLP-Large (1x) [18] 155.2 443.5 45.9 67.2 50.4 41.7 64.7 45.3
PVTv2-B5 [68] 101.6 334.5 47.4 68.6 51.9 42.5 65.7 46.0
ATMNet-L [69] 96.0 424.0 47.4 69.9 52.0 43.2 67.3 46.5
SpiralMLP-B (ours) 89.1 342.0 47.8 71.6 53.2 43.6 69.3 47.2
Table 3: Object detection performance with RetinaNet 1×1$\times$1 × and MASK R-CNN 1×1$\times$1 × on the COCO validation dataset, all of the backbones are pretrained on the ImageNet-1k. The FLOPs are evaluated at a resolution of $1280\times 800$. The entries are sorted in ascending order based on AP performance.
3.1.3 Comparison with other SOTAs
SpiralMLP remains competitive over Transformers, CNNs and State-Space Models, particularly in significantly reducing the number of parameters and the FLOPs as referenced in Tab. 2 , , . For instance, when comparing SpiralMLP-B5 to CNNs , it outperforms VanillaNet-13-1.5 [4] by +$1.5\%$ and has the same performance to DeepMAD-89M [51]. When comparing between State-Space Models and SpiralMLP-B4 as well as SpiralMLP-S, SpiralMLP demonstrates a notable performance improvement of approximate +$4.0\%$. Furthermore, when comparing with the Transformers , SpiralMLP-B5 has nearly $20$M fewer parameters than Swin-B [42] while achieving +$0.5\%$ higher in accuracy. Particularly the vision transformers continue struggling with quadratic complexity. And in order to better demonstrate, we visualize the heatmaps in Fig. 3 in comparison with the performance of ASMLP [36] and Swin [42].
3.2 Object Detection and Instance Segmentation on COCO
3.2.1 Settings
We conduct object detection and instance segmentation experiments on COCO [38], wherein we demonstrate SpiralMLP with PVT and Swin architectures, adopting two distinct configurations. We leverage SpiralMLP-B5 and Spiral-B with the pretrained weights on ImageNet-1k [50] as the backbones, together with Xavier initialization [16] applied to the newly added layers.
3.2.2 Results
Comparative results are detailed in Tab. 3, where we employ either RetinaNet [37] or Mask R-CNN [19] as the detection framework. When comparing under the RetinaNet 1×1$\times$1 ×, SpiralMLP-B5 stands out in terms of the highest AP. In particular, it achieves +$0.3\%$ higher than PVTv2-B5, with -$11.9$M fewer parameters. In the context of Mask R-CNN 1×1$\times$1 ×, SpiralMLP-B outperforms ATMNet-L by +$0.4\%$ in AP, alongside a reduction of $6.9$M in model parameters. Visual representations of the object detection and instance segmentation are presented in Fig. 4.
Figure 4: Several examples of the object detection and instance segmentation from COCO [38] test dataset.
Figure 5: Several examples of the semantic segmentation from ADE20K [82] validation dataset.
Model Semantic FPN Model UperNet
Params FLOPs mIoU Params FLOPs mIoU
ResNet101 [20] 47.5 10.1 38.8 DeepMAD-29M* [51] 27 56 46.9
ConvMLP-L [34] 46.3 - 40.0 HireMLP-Large [18] 127 1125 48.8
ResNeXt101-64x4d [71] 86.4 103.9 40.2 Focal-B [73] 126 - 49.0
CycleMLP-B5 [6] 79.4 86.0 45.5 ConvNeXt-T [43] 82 - 48.7
MorphMLP-B [79] 59.3 76.8 45.9 ConvNeXt-B [43] 122 - 49.1
Swin-B [42] 91.2 107.0 46.0 AS-MLP-B [36] 121 1166 49.5
Twins-L [8] 103.7 102.0 46.7 Swin-B [42] 121 1188 49.7
ConvNeXt-T [43] 27.8 93.2 46.7 CycleMLP-B [6] 121 1166 49.7
ATMNet-L [69] 79.8 86.6 48.1 ATMNet-L [69] 108 1106 50.1
PVTv2-B5 [68] 85.7 91.1 48.7 FocalNet-B(LRF) [72] 126 - 50.5
SpiralMLP-B5 (ours) 73.2 75.5 48.9 SpiralMLP-B (ours) 100 1061 50.7
Table 4: Semantic segmentation performance on ADE20K validation dataset with Semantic FPN as well as UperNet. When evaluated with Semantic FPN, the FLOPs are measured at a resolution of $512\times 512$. When evaluated with UperNet, the FLOPs are measured at a resolution of $2048\times 512$. The entries are sorted in ascending order based on mIOU performance.
3.3 Semantic Segmentation on ADE20K
3.3.1 Settings
We perform semantic segmentation on the ADE20K dataset using UperNet [70] and Semantic FPN [26] as the frameworks. For the backbones, we employ SpiralMLP-B5 and SpiralMLP-B, with the weights pretrained on ImageNet-1k. Additionally, the newly add layers are initialized with Xavier [16].
3.3.2 Results
As depicted in Tab. 4, SpiralMLP still exhibits comparable performance when integrated with Semantic FPN and UperNet for semantic segmentation tasks. In the Semantic FPN evaluations, SpiralMLP-B5 surpasses its closest competitor, PVTv2-B5, by +$0.2\%$, and exceeds the second-best model, ATMNet-L, by +$0.6\%$. When integrated with UperNet, SpiralMLP-B still emerges as the top-performing model, outperforming FocalNet-B(LRF) [72] by +$0.2\%$ and ATMNet-L [69] by +$0.6\%$. Visual representations of the semantic segmentation are presented in Fig. 5.
4 Ablation
Figure 6: Visualization of varying $k$ on spiral trajectory as described by Eqs. 14 and 15, while maintaining a constant $A_{\text{max}}=3$.
Case 1 $k$
$A_{\text{max}}=3$ 1 2 3 4 5
Acc(%) 83.9 84.0 83.8 83.6 83.3
Case 2 $A_{\text{max}}$
$k=2$ 2 3 4 5 6
Acc(%) 83.8 84.0 83.7 83.4 82.9
Table 5: Experiments on $k$ and $A_{\text{max}}$. After reaching their respective peaks, both trends show a rapid decline.
4.1 Update the Offset Functions
The offset functions $\phi_{i}(\cdot)$ and $\phi_{j}(\cdot)$ (Eqs. 4 and 5) are originally designed into a two-partition pattern, and we further expand them to a more generic multi-partition pattern. To incorporate this update, we introduce $k$ as the number of partitions along the channel dimension. The partitions can be defined as follows:
$P=\left\{0,\frac{C_{\text{in}}}{k},\frac{2*C_{\text{in}}}{k},\dots,C_{\text{in}}\right\}$ (12)
By introducing $k$ and considering individual partition, we can create multiple spiral structures that capture the characteristics of each partition along the channel dimension. Furthermore, we define the length of the partition as $C_{w}=\frac{C_{\text{in}}}{k}$, which is the distance between two adjacent endpoints, then the amplitude function Eq. 6 is updated to:
$A^{*}(c)=\begin{cases}\left\lfloor\frac{2A_{\text{max}}}{C_{w}}(c-iC_{w})\right\rfloor,&0\leq c<\frac{iC_{w}}{2}\\ \left\lfloor(2A_{\text{max}}-\frac{2A_{\text{max}}}{C_{w}})(c-iC_{w})\right\rfloor,&\frac{iC_{w}}{2}\leq c\leq iC_{w}\end{cases}$ (13)
where, $i\in[0,1\dots,k-1]$ represents the $i^{\text{th}}$ partition in partitions $P$ (Eq. 12), and $c$ in Eq. 6 is replaced by $z$ within the $i^{\text{th}}$ partition. Accordingly, Eqs. 4 and 5 are updated as:
$\phi^{*}_{i}(c)$ $=A^{*}(c)\cos(\frac{c*2\pi}{T})$ (14)
$\phi^{*}_{j}(c)$ $=A^{*}(c)\sin(\frac{c*2\pi}{T})$ (15)
We also provide the visualizations of Eqs. 14 and 15, as depicted in Fig. 6, showcasing variations with different numbers of partitions $k$.
SpiralFC PATM ATMLayer CycleFC RandomFC
Acc (%) 95.6 95.3 95.2 94.7 94.5
Params (M) 14 17 15 15 14
Table 6: The accuracy on CIFAR-10, each Fully-Connected Layer is configured into the SpiralMLP-B1 architecture and is trained from scratch.
4.2 Ablation Study on $k$
We updated the offset functions $\phi^{*}_{i}(\cdot)$ and $\phi_{j}^{*}(\cdot)$ (Eqs. 14 and 15) to analyze how varying the number of partitions $k$ affects Top-1 Accuracy on ImageNet-1k. Results, shown in Tab. 5 with a constant maximum amplitude $A_{\text{max}}$ of $3$, indicate that accuracy initially rises, peaks at $k=2$, then decreases.
This trend suggests that different $k$ values alter the focus on the peripheral regions of the receptive field, where $k=2$ results in denser clustering of feature points along the edges compared to $k=4$, as seen in Fig. 6. Lower $k$ values cause a dense, narrow concentration of features, while higher values disperse them too widely, potentially reducing model effectiveness.
4.3 Ablation Study on $A_{\text{max}}$
We investigate several cases when the maximum amplitude $A_{\text{max}}$ takes various values. From the results shown in Tab. 5, we observe an initial improvement in the Top-1 Accuracy on ImageNet-1k. However, a decline becomes evident once the $A_{\text{max}}$ exceeds $3$.
Similarly, the underlying reason is that, as $A_{\text{max}}$ increases, the receptive field’s extent expands. However, due to the characteristics of Spiral FC, the number of selected feature points remains constant at $C_{\text{in}}$. Consequently, a larger $A_{\text{max}}$ results in a more sparse distribution of feature points. If $A_{\text{max}}$ is too small, the Spiral FC may fail to encompass a sufficient number of neighboring features. On the other hand, if $A_{\text{max}}$ is excessively large, the Spiral FC might not effectively capture detailed information within the receptive field.
Although the discrete experimental design does not guarantee the discovery of optimal hyperparameters, it indeed facilitates the insight of underlying trends and tendencies.
4.4 Ablation Study on Fully-Connected Layers
To illustrate the effectiveness of Spiral FC, we perform experiments on the CIFAR-10 [29] using SpiralMLP-B1333The configuration of SpiralMLP-B1 is demonstrated in Appendix. as the base architecture. In these experiments, the Spiral FC is substituted with various alternatives, including PATM from WaveMLP [58], ATMLayer from ATM [69], CycleFC from CycleMLP [6] and a RandomFC. The Random FC is architecturally identical to Spiral FC, except that the offset function is generated randomly.
4.5 Latency Analysis
To highlight the speed efficiency of Spiral FC, we assess its performance across various input resolutions compared to other proposed architectures. We adopt the format from EfficientFormer [35] and detail the latency analysis in  Tab. 7. We present SpiralMLP-B4 and SpiralMLP-B5 with several other architectures closely related to our study and specifically at the $224^{2}$ resolution. For a comprehensive latency comparison across different scenarios, please refer to the Appendix.
Model Params(M) Latency(ms)
SpiralMLP-B4 46 47.00
SpiralMLP-B5 68 39.22
CycleMLP-B4 [6] 52 57.94
CycleMLP-B5 76 48.38
ATM-B [69] 52 64.77
ATM-L 76 54.09
PVTv2-B4 [68] 63 43.96
PVTv2-B5 82 55.71
Table 7: Inference latency measured in milliseconds on one A100. SpiralMLP outperforms other models of similar size in terms of speed. A single image of with $224^{2}$ resolution serves as the input.
5 Conclusion and Future Work
In this paper, we present Spiral FC, part of Spiral Mixing designed to replace traditional Token Mixing. We introduce SpiralMLP, a new computer vision framework compatible with PVT-style and Swin-style architectures. SpiralMLP performs comparably to leading models while using fewer parameters and less computational power.
We believe we are the first to use carefully designed offset functions to capture comprehensive feature information, setting us apart from models like CyelcMLP [6], ASMLP [36] and ATM [69], which focus on optimizing cross-like layers. Given its strong performance, further research into optimizing SpiralMLP’s hyperparameters could lead to even more efficient information capture.
Appendix A Related Works
A.1 CNN-Based
For a long range of time, CNN-based architectures have dominated the computer vision domain. The prototype of CNN is presented in [31] and after the exciting success of AlexNet [30], a large number of methods adopt CNN architecture for higher performance [52, 55, 20]. Especially the ResNet [20], utilizes a residual connection among layers to alleviate the gradient vanishing. Owing to the hierarchical structure employed in CNNs, rich information can be effectively extracted from localized receptive fields. Yet, despite advantages, CNNs exhibit certain limitations, particularly in terms of capturing global contextual information, and inductive bias both could potentially impede CNNs being applied to downstream tasks.
A.2 Transformer-Based
The foundation work [66] on Transformer-based architecture proposes the attention mechanism used to extract the relationship among the spatial positional features. Following the groundbreaking achievements of BERT [12] in natural language processing, numerous approaches have leveraged Transformer-based architecture for advanced performance [40, 54, 49, 74, 48, 3, 45, 63, 64, 14, 2, 42], both in natural language processing and computer vision. Notably with self-attention mechanism, Vision Transformer [14] has facilitated long-range dependencies, indicating that Transformers-based architectures have the ability to extract global context understanding.
Despite remarkable achievements across various domains, Transformers do have certain drawbacks that need to be considered. $\text{SoftMax}(\frac{Q\times K^{T}}{\sqrt{k}})V$, is the primary reason for heavy computational demands. Furthermore, the scalability is not only impeded by its quadratic computational complexity, but also by the necessity for extensive datasets and significant memory consumption.
Appendix B Compare SpiralMLP to CNNs and MHSA
In order to establish a comparison between Multi-Head Self-Attention (abbreviated as MHSA) and SpiralMLP, it is necessary to demonstrate the relationship between MHSA and Convolutional Neural Networks (abbreviated as CNNs). This is because SpiralMLP, unlike MHSA, does not incorporate self-attention layers and is more closely aligned with CNNs.
By examining the connection between MHSA and CNNs, we can provide a comprehensive understanding of the architectural differences and similarities between MHSA and SpiralMLP. This will enable us to highlight the unique features and advantages of each approach.
Thus we outline the following subsections in this way:
  • We first elaborate how SpiralMLP is related to CNNs.
  • Then we draw a relation between MHSA and CNNs using the proof of Cordonnier [9].
  • Finally we provide the comparison between MHSA and SpiralMLP.
B.1 How is SpiralMLP related to CNNs?
In order to demonstrate the functionality of CNNs, we define the standard convolution weight matrix as $W^{\text{cnn}}\in\mathbb{R}^{K_{H}\times K_{W}\times C_{\text{in}}\times C_{\text{out}}}$, where $K_{H}$, $K_{W}$, $C_{\text{in}}$, $C_{\text{out}}$ are the kernel height, kernel width, input channel dimension, output channel dimension, respectively. Given a position $(i,j,:)$ within the feature map $X\in\mathbb{R}^{H\times W\times C_{\text{in}}}$, where $H$, $W$ are the height and weight, and $\Omega=\{(x,y);0\leq x\leq K_{W},0\leq y\leq K_{H}\}$ is defined as a set of coordinates in a rectangular, corresponding to the kernel. Thus the convolution operation is formulated as:
$\text{Conv}_{i,j,:}(X)=\sum_{(x,y)\in\Omega}X_{i+x,j+y,:}W^{\text{cnn}}_{y,x,:,:}+b^{\text{cnn}}$ (16)
where, $b^{\text{cnn}}\in\mathbb{R}^{C_{\text{out}}}$ is the CNNs bias, the overall output $\text{Conv}_{i,j,\colon}(\cdot)$ is a vector at each position $(i,j,:)$.
Next, we replace the dot product between $X_{i+x,j+y,:}$ and $W^{\text{cnn}}_{y,x,:,:}$ in Eq. 16 with a summation across the channel dimension to get:
$\text{Conv}_{i,j,:}(X)=\sum_{(x,y)\in\Omega}\sum^{C_{\text{in}}}_{c=0}X_{i+x,j+y,c}W^{\text{cnn}}_{y,x,c,:}+b^{\text{cnn}}$ (17)
Afterwards we change the order of the summation and obtain:
$\text{Conv}_{i,j,:}(X)=\sum^{C_{\text{in}}}_{c=0}\sum_{(x,y)\in\Omega}X_{i+x,j+y,c}W^{\text{cnn}}_{y,x,c,:}+b^{\text{cnn}}$ (18)
To relate Eq. 18 to the Spiral FC (Eqs. 3, 4 and 5), we first create a function:
$g(x,y,c)=\begin{cases}1,\quad&\text{if}\ x=\phi_{i}(c)\ \text{and}\ y=\phi_{j}(c)\\ 0,\quad&\text{otherwise}\end{cases}$ (19)
Then we apply the Eq. 19 on the kernel weights $W^{\text{cnn}}$ of the CNNs to obtain:
$\widetilde{W}^{\text{cnn}}_{y,x,c,:}=g(x,y,c)\cdot W^{\text{cnn}}_{y,x,c,:}$ (20)
Subsequently, we substitute the weights defined in Eq. 18 with the weights in Eq. 20:
$\text{Conv}_{i,j,:}(X)$ (21)
$=\sum^{C_{\text{in}}}_{c=0}\sum_{(x,y)\in\Omega}X_{i+x,j+y,c}\widetilde{W}^{\text{cnn}}_{y,x,c,:}+b^{\text{cnn}}$
$=\sum^{C_{\text{in}}}_{c=0}\sum_{(x,y)\in\Omega}g(x,y,c)X_{i+x,j+y,c}W^{\text{cnn}}_{y,x,c,:}+b^{\text{cnn}}$
$=\sum^{C_{\text{in}}}_{c=0}X_{i+\phi_{i}(c),j+\phi_{j}(c),c}W^{\text{spiral}}_{c,:}+b^{\text{spiral}}$
$=\text{Spiral FC}_{i,j,:}(X)$
where, $W^{\text{spiral}}$ and $b^{\text{spiral}}$ are already introduced in Eq. 3.
Consequently, we have demonstrated how the Spiral FC can be derived from convolutions. It is noteworthy that Eqs. 19, 20 and 21 reveal that the Spiral FC assumes sparsity within the convolution operation, indicating its potential advantages and distinctive characteristics compared to traditional dense convolution layers.
B.2 How is MHSA related to CNNs?
To formulate, the multi-head self-attention weight matrix for each head is denoted as $W^{\text{mhsa,}h}\in\mathbb{R}^{C_{\text{in}}\times C_{\text{out}}}$, and the bias is denoted as $b^{\text{mhsa}}\in\mathbb{R}^{C_{\text{out}}}$. Furthermore, $h$ refers to each individual head and $N_{h}$ represents the total number of heads. The process of MHSA applied to the input $X\in\mathbb{R}^{H\times W\times C_{\text{in}}}$ on position $(i,j,:)$ can be arranged as:
$\text{MHSA}_{i,j,:}(X)=\sum^{N_{h}}_{h=1}X_{:,:,C^{h}_{\text{in}}}W^{\text{mhsa},h}_{C^{h}_{\text{in}},:}+b^{\text{mhsa}}$ (22)
where, $C^{h}_{\text{in}}$ illustrates the division of the input channel corresponding to the $h^{\text{th}}$ head, and it is incorporated into the computation for each head $h$.
While in the work [9], it is demonstrated that MHSA can be represented in a manner similar to CNNs:
$\text{MHSA}_{i,j,:}(X)$ (23)
$=\sum^{N_{h}}_{h=1}X_{i+\Delta_{i}(h),j+\Delta_{j}(h),C^{h}_{\text{in}}}W^{\text{mhsa},h}_{C^{h}_{\text{in}},:}+b^{\text{mhsa}}$
$=\sum^{N_{h}}_{h=1}\sum^{C^{h}_{\text{in}}}_{c=0}X_{i+\Delta_{i}(h),j+\Delta_{j}(h),c}W^{\text{mhsa},h}_{c,:}+b^{\text{mhsa}}$
where, all relative positional shifts for a kernel of size $\sqrt{N_{h}}\times\sqrt{N_{h}}$ at position $(i,j)$ are contained in $\{\Delta_{i}(h),\Delta_{j}(h)\}=\{(-1,0),(-1,1),(-1,2),\dots\}$.
B.3 How is SpiralMLP related to MHSA?
Through a similar technique to Eq. 20, we establish a connection between MHSA and SpiralMLP through CNNs. Firstly, we augment $W^{\text{mhsa},h}$ by multiplying with the Eq. 19 to obtain:
$\widetilde{W}^{\text{mhsa},h}_{c,:}=g(x,y,c)\cdot W^{\text{mhsa},h}_{c,:}$ (24)
where, $\widetilde{W}^{\text{mhsa},h}$ is dependent on the position along the kernel height, width and input channel. Substitute it into Eq. 23:
$\text{MHSA}_{i,j,:}(X)$ (25)
$=\sum^{N_{h}}_{h=1}\sum^{C^{h}_{\text{in}}}_{c=0}X_{i+\Delta_{i}(h),j+\Delta_{j}(h),c}\widetilde{W}^{\text{mhsa},h}_{c,:}+b^{\text{mhsa}}$
$=\sum^{N_{h}}_{h=1}\sum^{C^{h}_{\text{in}}}_{c=0}X_{i+\phi_{i}(c),j+\phi_{j}(c),c}W^{\text{mhsa},h}_{c,:}+b^{\text{mhsa}}$
$=\sum^{C_{\text{in}}}_{c=0}X_{i+\phi_{i}(c),j+\phi_{j}(c),c}W^{\text{spiral}}_{c,:}+b^{\text{spiral}}$
$=\text{Spiral FC}_{i,j,:}(X)$
where, $W^{\text{spiral}}$ and $b^{\text{spiral}}$ are already introduced in Eq. 3.
Eq. 25 demonstrates the relationship between Spiral FC and MHSA. Drawing a parallel with the relationship between CNNs and Spiral FC, we can conclude that Spiral FC exhibits a significantly sparser receptive field compared to MHSA. This highlights the distinctive characteristic of Spiral FC in terms of its sparse attention mechanism compared to the dense attention mechanism of MHSA.
Output Size Layer Name PVT-Style
B1 B2 B3 B4 B5
Stage1 $\frac{H}{4}\times\frac{W}{4}$ Overlapping $S_{1}=4$
Patch Embedding $C_{1}=64$ $C_{1}=96$
SpiralMLP $E_{1}=4$ $E_{1}=4$ $E_{1}=4$ $E_{1}=4$ $E_{1}=4$
Block $L_{1}=2$ $L_{1}=2$ $L_{1}=3$ $L_{1}=3$ $L_{1}=3$
Stage2 $\frac{H}{8}\times\frac{W}{8}$ Overlapping $S_{2}=2$
Patch Embedding $C_{2}=128$ $C_{2}=192$
SpiralMLP $E_{2}=4$ $E_{2}=4$ $E_{2}=4$ $E_{2}=4$ $E_{2}=4$
Block $L_{2}=2$ $L_{2}=3$ $L_{2}=4$ $L_{2}=8$ $L_{2}=4$
Stage3 $\frac{H}{16}\times\frac{W}{16}$ Overlapping $S_{3}=2$
Patch Embedding $C_{3}=320$ $C_{3}=384$
SpiralMLP $E_{3}=4$ $E_{3}=4$ $E_{3}=4$ $E_{3}=4$ $E_{3}=4$
Block $L_{3}=4$ $L_{3}=10$ $L_{3}=18$ $L_{3}=27$ $L_{3}=24$
Stage4 $\frac{H}{32}\times\frac{W}{32}$ Overlapping $S_{4}=2$
Patch Embedding $C_{4}=512$ $C_{4}=768$
SpiralMLP $E_{4}=4$ $E_{4}=4$ $E_{4}=4$ $E_{4}=4$ $E_{4}=4$
Block $L_{4}=2$ $L_{4}=3$ $L_{4}=3$ $L_{4}=3$ $L_{4}=3$
Parameters (M) 14 24 34 46 68
FLOPs (G) 2.0 3.6 5.6 8.2 11.3
Accuracy Top-1 (%) 79.8 81.9 83.4 83.8 84.0
Table 8: Configurations of SpiralMLP variants in PVT-style.
Appendix C Model Zoo and Training Details
C.1 Model Zoo Configurations and Performances
We implement five models according to PVT style, named SpiralMLP B1 to B5, and three models based on Swin-style, named as SpiralMLP-T, SpiralMLP-S, SpiralMLP-B.
SpiralMLP variants in PVT-style are detailed in Tab. 8. Each stage contains multiple Spiral Blocks with uniform configurations, where the parameters $S$, $E$, $C$, and $L$ denote the shift size, expansion ratio, channel dimension, and the number of layers in each stage, respectively.
SpiralMLP variants in Swin-style are shown in Tabs. 9 and 7. The input image dimension is $224\times 224$. The process ’concat $n\times n$’ refers to the concatenation of features from $n\times n$ neighboring features in a patch. This operation effectively downsamples the feature map by a factor of $n$. The notation ’96-d’ represents a linear layer whose output dimension is 96. In the term [(3,2), 96] indicates that $A_{\text{max}}=3$, $k=2$, $C_{\text{out}}=96$.
C.2 Experimental Setup for Image Classification
[…]
C.3 Experimental Setup for Object detection and Instance Segmentation
[…]
C.4 Experimental Setup for Semantic Segmentation
The semantic segmentation is conducted on ADE20K [82], which consists of 20k training images and 2k validation images. The frameworks used for this purpose are Semantic FPN [26] and UperNet [70], employing SpiralMLP with ImageNet-1k pretrained weights as the backbones. For optimization, the AdamW [44] optimizer is chosen. In the case of Mask R-CNN, the optimizer is set with an initial learning rate of $0.0001$ and a weight decay of $0.05$. For Semantic FPN, the same initial learning rate of $0.0001$ is used, but the weight decay is lower, at $0.0001$. The training process is carried out on $4$ A100 GPUs. A batch size of $32$ is maintained throughout, and the model undergoes training for $12$ epochs.
Output Size SpiralMLP-T SpiralMLP-S SpiralMLP-B
Stage1 $4\times(56\times 56)$ concat $4\times 4$, 64-d, LN concat $4\times 4$, 96-d, LN concat $4\times 4$, 96-d, LN
[(3,2), 64]$\times$2 [(3,2), 96]$\times$3 [(3,2), 96]$\times$3
Stage2 $8\times(28\times 28)$ concat $2\times 2$, 128-d, LN concat $2\times 2$, 192-d, LN concat $2\times 2$, 192-d, LN
[(3,2), 128]$\times$2 [(3,2), 192]$\times$4 [(3,2), 192]$\times$4
Stage3 $16\times(14\times 14)$ concat $2\times 2$, 320-d, LN concat $2\times 2$, 384-d, LN concat $2\times 2$, 384-d, LN
[(3,2), 320]$\times$6 [(3,2), 384]$\times$18 [(3,2), 384]$\times$24
Stage4 $32\times(7\times 7)$ concat $2\times 2$, 512-d, LN concat $2\times 2$, 768-d, LN concat $2\times 2$, 768-d, LN
[(3,2), 512]$\times$2 [(3,2), 768]$\times$3 [(3,2), 768]$\times$3
Parameters (M) 15 56 67
FLOPs (G) 2.3 9.1 11.0
Accuracy Top-1 (%) 79.6 83.3 83.6
Table 9: Configurations of SpiralMLP variants in Swin-style.
Figure 7: The architecture of SpiralMLP in Swin-style.
Appendix D Merge Head
D.1 How does the Merge Head Work?
In Eq. 7, Merge Head employs a trainable weights $a$ to determine the contribution of $X^{\text{self}}\in\mathbb{R}^{H\times W\times C_{\text{out}}}$, $X^{\text{cross}}\in\mathbb{R}^{H\times W\times C_{\text{out}}}$, from Channel FC and Spiral FC, respectively. For a clearer understanding, we present the tensor shape of each step in Merge Head, as shown in Tab. 10. Rather than merely employing a straightforward addition, the merge head innovatively takes into account the inputs themselves to formulate the trainable weights $a$. This approach allows for a more dynamic and input-responsive weight adjustment, enhancing the effectiveness of the merging process.
D.2 Complexity of Merge Head.
Two inputs are both of shape $\mathbb{R}^{H\times W\times C}$. The addition has $O(HW)$. While the reshaping operation $\mathcal{F}(\cdot)$ is implemented by torch.flatten() with $O(1)$. The average calculation is implemented by torch.mean() with $O(HW)$. The multiplication involves $W^{\text{merge}}\in\mathbb{R}^{2,1}$ contributes $O(1)$. Therefore, the total complexity of this process is linear $O(HW)$.
Step Tensor Shape
$(X^{\text{self}},X^{\text{cross}})$ $(\mathbb{R}^{H\times W\times C_{\text{out}}},\mathbb{R}^{H\times W\times C_{\text{out}}})$
$\mathcal{F}(\cdot)$ $\mathbb{R}^{H\cdot W\times C_{\text{out}}}$
Average $\mathbb{R}^{1\times C_{\text{out}}}$
$W^{\text{merge}}$ $\mathbb{R}^{2\times C_{\text{out}}}$
Table 10: Tensor shape of each step in Merge Head.
Appendix E Why Spiral FC works?
Eq. 21 indicates that SpiralMLP, along with other criss-cross MLPs, can be effectively implemented using a specialized convolution layer. Currently, deformable convolution [11] emerges as the optimal method for this purpose.
In the given scenario as shown in Fig. 8, points are identified as $(o_{x},o_{y})$, $(\mu_{x},\mu_{y})$, $(\nu_{x},\nu_{y})$, and $(\mu_{x},\nu_{y})$, all located on feature map $X$. Specifically, when $(o_{x},o_{y})$ is considered the reference point, the points $(\mu_{x},\mu_{y})$ and $(\nu_{x},\nu_{y})$ are categorized into set $P$. This set includes pairs that are either horizontally or vertically aligned with the reference point $(o_{x},o_{y})$. The point $(\mu_{x},\nu_{y})$ falls into set $Q$, which comprises pairs that can be located at any position within the feature map $X$. Then the set $P$ and set $Q$ are defined as:
Figure 8: Spiral FC covers a more comprehensive receptive field.
$P$ =$=\{(\mu_{x},\mu_{y})\in\mathbb{R}^{2}|(\mu_{x}=o_{x}\lor\mu_{y}=o_{y})\}$ (26)
$Q$ =$=\{(\mu_{x},\nu_{y})\in\mathbb{R}^{2}\}$ (27)
We further denote the $Y_{C}\in\mathbb{R}^{H\times W\times C}$ as the output of the convolution layer, while $X\in\mathbb{R}^{H\times W\times C}$ is still the feature map, this is similar to Eq. 21:
$Y_{C}=\sum^{C_{\text{in}}}_{c=0}X_{i+\bar{\phi}_{i}(c),j+\bar{\phi}_{j}(c),c}W_{c,:}+b$ (28)
where, $\bar{\phi}_{i}(\cdot),\bar{\phi}_{j}(\cdot)$ denote the universal offset functions.
We take two architectures with criss-cross offset functions as examples. When using CyelcMLP [6] as an example, the offset function is updated into:
$\phi^{\text{cycle}}_{i}(c)$ $=(c\mod S_{H})-1$ (29)
$\phi^{\text{cycle}}_{j}(c)$ $=(\lfloor\frac{c}{S_{H}}\rfloor\mod S_{W})-1$ (30)
where, $S_{H},S_{W}$ are the predefined step size.
When using ASMLP [36], the offset function is updated into:
$\phi^{\text{as}}_{i}(c)$ $=\lfloor\frac{c}{C/s}\rfloor-\lfloor\frac{s}{2}\rfloor\cdot d$ (31)
$\phi^{\text{as}}_{j}(c)$ $=\lfloor\frac{c}{C/s}\rfloor-\lfloor\frac{s}{2}\rfloor\cdot d$ (32)
where, $s$ is the shift size and $d$ is the dilation rate.
In Fig. 8 (a), points $(\mu_{x},\mu_{y})$ and $(\nu_{x},\nu_{y})$ from set $P$ are reachable by the offset functions $\phi^{\text{cycle}}(\cdot)$, and $\phi^{\text{as}}(\cdot)$ due to their placement along the horizontal or vertical axes. However, as illustrated in Fig. 8 (b), points that do not lie on these axes pose a challenge for the criss-cross methodology, which inherently lacks the capability to capture such spatial information effectively.
To address this limitation, Fig. 8 (c) suggests the adoption of a predefined multi-helix offset approach. This method, while effective, still offers room for optimization to achieve model efficiency in terms of size and computational speed. A viable solution is the refinement of the multi-helix offset into a spiral-like offset function. This adjustment not only enables the model to recognize points from set $Q$—those not aligned with the horizontal or vertical axes—but also maintains a compact model architecture and ensures rapid processing speeds.
Model Params(M) Throughput (/sec) Wall-Clock (s)
Spiral-B1 14 425 22.8
Spiral-B2 24 253 30.5
Spiral-B3 34 143 52.2
Spiral-B4 46 94 72.5
Spiral-B5 68 89 79.2
Cycle-B1 [6] 15 347 23.8
Cycle-B2 27 207 36.0
Cycle-B3 38 117 60.4
Cycle-B4 52 79 84.0
Cycle-B5 76 75 92.0
ResNet-152 [20] 60 152 45.7
Vit-B/16 [14] 86 102 75.2
Deit-B/16 [61] 86 97 78.3
Table 11: Throughput and wall-clock test.
Appendix F Latency & Throughput Analysis
As shown in Tab. 11, throughput testing is conducted on a single NVIDIA A100 GPU, with a batch size of 32 and image resolution of 3x224x224; wall-clock testing is conducted on ImageNet-1k with a single NVIDIA A100 GPU, with a batch size of 64, the number of total batches to 100 and image resolution of 3x224x224. Furthermore, we evaluate the latency speeds on one A100 across various architectures for input resolutions of $224^{2}$, $384^{2}$, and $512^{2}$. These results are detailed in  Tab. 12.
It is evident that SpiralMLP is faster than other MLP-based models when considering the scale of model size. Additionally, the PVT [67] showcases even faster processing speeds, mainly due to its utilization of dot product operations, inherently benefiting from accelerated computation. The models used for evaluation are directly extracted from their official implementations.
Model $224^{2}$ $384^{2}$ $512^{2}$ Params(M) Model $224^{2}$ $384^{2}$ $512^{2}$ Params(M)
Spiral-B1 11.74 11.83 11.52 14 Cycle-B1 [6] 12.12 12.06 12.23 15
Spiral-B2 20.81 21.11 20.29 24 Cycle-B2 21.66 21.36 21.12 27
Spiral-B3 32.33 32.53 31.38 34 Cycle-B3 33.23 32.89 32.29 38
Spiral-B4 47.00 47.39 45.95 46 Cycle-B4 47.86 48.29 46.73 52
Spiral-B5 39.22 39.34 38.34 68 Cycle-B5 41.00 40.87 39.59 76
ATM-xT [69] 15.01 15.32 14.86 15 Wave-T-dw [58] 16.91 15.83 16.01 15
ATM-T 25.83 26.66 25.43 27 Wave-T 18.56 17.48 17.47 17
ATM-S 37.47 38.41 37.21 39 Wave-S 31.97 30.76 30.61 31
ATM-B 55.45 55.85 53.39 52 Wave-M 48.98 47.43 46.94 44
ATM-L 47.24 45.98 44.78 76 Wave-B 41.99 40.50 40.27 64
PVT-Tiny [67] 8.54 8.72 8.55 13 PVTv2-B1 [68] 9.08 9.04 9.12 14
PVT-Small 15.73 15.75 15.62 25 PVTv2-B2 17.14 17.45 17.21 25
PVT-Medium 26.78 27.48 26.71 44 PVTv2-B3 30.57 30.21 30.32 45
PVT-Large 39.28 39.75 38.68 61 PVTv2-B4 44.77 44.73 43.51 63
- - - - - PVTv2-B5 56.64 56.88 55.19 82
Table 12: Latency analysis measured in milliseconds(ms) on one A100. A single image with differing resolutions serves as the input
Appendix G Resolution Compatibility
We compare the image classification models resolution compatibility with Top-1 accuracy on ImageNet-1k, with models pre-trained on 224 x 224 images and tested at various other resolutions without additional fine-tuning, as shown in Tab. 13.
Model Params(M) 128 224 256 384
Spiral-B5 68 77.6 84.0 83.8 83.4
Cycle-B [6] 88 77.0 83.2 83.1 82.7
ViT-B/16 [14] 86 - 81.0 81.6 79.0
ResNet-152 [20] 60 68.7 78.3 78.4 77.4
DeiT-B/16 [61] 87 - 81.7 - 82.5
Table 13: Resolution compatibility comparison.
Appendix H Spiral in AttentionViz
In the work AttentionViz [75], the authors introduce a visualization tool for analyzing transformer models, particularly focusing on the interaction between keys and queries distributions across various heads and layers. A notable discovery in their research is the identification of a spiral-like pattern. This pattern suggests that the keys and queries in transformers are spatially aligned in a manner resembling a spiral. The phenomenon results from position vectors that are generated using trigonometric functions, mapping onto a helical curve in a high-dimensional space. In linguistics models, this refers to the arrangement of words or parts of words; while in vision transformers, it relates to the organization of pixel patches.
This spiral distribution is reflective of the initial ordering vector given to transformers, illustrating how positional information is embedded within the model. Furthermore, the research reveals that transformer heads displaying a spiral shape or having clusters of queries/keys tend to yield more dispersed search results.